Calculating the Predetermined Overhead Rate, Applying Overhead to Production, Reconciling Overhead at the End of the Year, Adjusting Cost of Goods Sold for Under- and Overapplied Overhead

At the beginning of the year, Horvath Company estimated the following:

Overhead $360,000
Direct labor hours 90,000

Horvath uses normal costing and applies overhead on the basis of direct labor hours. For the month of January, direct labor hours were 8,550. By the end of the year, Horvath showed the following actual amounts:

Overhead $366,000
Direct labor hours 89,600

Assume that unadjusted Cost of Goods Sold for Horvath was $576,000.

Required:

1. Calculate the predetermined overhead rate for Horvath. Round your answers to the nearest cent, if rounding is required.
$ per direct labor hour

2. Calculate the overhead applied to production in January. (Note: Round to the nearest dollar, if rounding is required.)
$

3. Calculate the total applied overhead for the year.
$

Was overhead over- or underapplied? By how much?
  overhead $

4. Calculate adjusted Cost of Goods Sold after adjusting for the overhead variance.
$
Answer #1

1.

Estimated overhead = $360,000

Estimated direct labor hours = 90,000

Predetermined overhead rate = Estimated overhead / Estimated direct labor hours

= 360,000/90,000

= $4 per direct labor hour

2.

Actual direct labor hour used = 8,550

Overhead applied in January = Actual direct labor hour used x Predetermined overhead rate

= 8,550 x 4

= $34,200

3.

Actual direct labor hours used in the year = 89,600

Overhead applied for the year = Actual direct labor hours used in the year x Predetermined overhead rate

= 89,600 x 4

= $358,400

Actual overhead = $366,000

Under Applied overhead = Actual overhead - Applied overhead

= 366,000-358,400

= $7,600

Since applied overhead was less than actual overhead. Hence, overhead was under applied.

4.

Unadjusted Cost of goods sold = $576,000

Adjusted cost of goods sold = Unadjusted cost of goods sold + Under applied overhead

= 576,000+7,600

= $583,600

Please give a positive rating if you are satisfied with this solution and if you have any query kindly ask.

Thanks!!!